OpenStudy (anonymous):

Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately what maximum height does the soccer ball reach?

OpenStudy (anonymous):

at wht angle to the ground is the ball kicked?

OpenStudy (amistre64):

the angle doesnt matter..... just gotta use the gravity formula... -16t o something

OpenStudy (amistre64):

asuume straight if anything :)

OpenStudy (amistre64):

striaght up... that is

OpenStudy (anonymous):

v=gt t=v/g=4.2seconds s=1/2gt^2=5*4.2^2=88.2feet

myininaya (myininaya):

http://www.purplemath.com/modules/quadprob.htm i think this site is helpful for writing formula for velocity

myininaya (myininaya):

or i mean height of the ball

myininaya (myininaya):

s(t)=-16t^2+42t+0

OpenStudy (amistre64):

s(t) = –G t^2 + V0(t) + H0 Vo = 42 ; H0 = 0; G = 16 s(t) = -16t^2 + 42t s'(t) = -32t + 42 = 0 at 42/32 = 21/16 = 1.3125

myininaya (myininaya):

s'(t)=-32t+42 s'(t)=0 when -32t+42=0 so t=42/32=21/16

myininaya (myininaya):

lol

OpenStudy (amistre64):

s(t) = -16(1.3125^2) + 42(1.3125) = 27.5625 :) maybe

myininaya (myininaya):

k i agree with amistre

OpenStudy (amistre64):

yay!!.... if great minds think alike; what does that leave for idiots lol

myininaya (myininaya):

we got something way different from that other guy

OpenStudy (amistre64):

he mighta used the 4.9 meters thing

myininaya (myininaya):

ok in this is in feet

myininaya (myininaya):

so maybe

OpenStudy (anonymous):

hello all. equation is \[42t-16t^2\] max is when \[t=\frac{-b}{2a}=\frac{21}{16}\] \[42\times \frac{21}{16} -16\times(\frac{21}{16})^2=\frac{441}{16}\]

myininaya (myininaya):

we already win :)

OpenStudy (anonymous):

;)

OpenStudy (amistre64):

yay!!! now if we can only determine that we are not a consensus of fools; I for one have my doubts since I am in it lol

OpenStudy (anonymous):

btw you have to use \[h=h_0+v_0t-16t^2\] because you are working in feet. and you don't need calc to find the vertex of a parabola, although you can

OpenStudy (amistre64):

i did just casue I can :)

myininaya (myininaya):

right satelitle! derivatives are just more fun!

OpenStudy (anonymous):

i spose but after getting \[\frac{-b}{2a}\]for the 100th time maybe not.

myininaya (myininaya):

i like doing more math than i need to because its just fun lol

OpenStudy (amistre64):

more fun; less mechanical ;)

OpenStudy (anonymous):

on the other hand sometime it is silly to kill a fly with a grenade. you can get the vertex by simple algebra, completing the square or something.

myininaya (myininaya):

omg i would love to kill a fly with a grenade how is that not fun

OpenStudy (anonymous):

so not to be argumentative but i would not consider taking the derivative, setting to 0 and solving more or less mechanical than computing the vertex.

myininaya (myininaya):

lol ok you do that :)

myininaya (myininaya):

i might do it both ways just to show both ways and to do more MATH

OpenStudy (anonymous):

what i need to do is get back to my own work but somehow i keep coming here.

OpenStudy (amistre64):

i know right; it worse than cocaine in here lol